Caculate the volume, in ml of 0.2 M NaOH solution that contains 6.0 g of NaOH.

1
Expert's answer
2022-04-13T15:20:01-0400

0.2M means 0.2 mole/l

M(NaOH)=23+16+1=40 g/mole

n(NaOH)=6/40=0.15 mole

"0.2=\frac{0.15}{V}"

V=0.15/0.2=0.75 l=750 ml